OpenStudy (anonymous):

Amy has a triangular flag. She wants to cut out a similar flag that has dimensions one-sixth of the original flag. If the area of the original flag is 2160 square inches, what is the area of the new flag that Amy cuts out? Answer 10 in2 20 in2 50 in2 60 in2

OpenStudy (anonymous):

Why do you delete your earlier question?

OpenStudy (sasogeek):

I believe it's 60... but please don't delete your earlier questions :) people might need it for reference :)

OpenStudy (anonymous):

Arent you suppose to after you get answer?

OpenStudy (anonymous):

oh ok sorry :(

OpenStudy (sasogeek):

nope :)

OpenStudy (sasogeek):

it's all good :) i believe you're fairly new here so just learn how things work around here, you'll get used to it :)

OpenStudy (anonymous):

So you all think the answer is 60?

OpenStudy (anonymous):

saso, explain your thinking ;-)

OpenStudy (sasogeek):

lol well here's how i arrived at 60. 1/6 of 2160 = 360 and for some reason i divided by 6 and had 60. i'm trying to figure out why i divided by 6 but nothing is coming, maybe my brain was thinking ahead of me... ffm you might want to explain this mathematically...

OpenStudy (anonymous):

mathematically ? I can only guess :P

OpenStudy (sasogeek):

well 1/6 of 2160 is 360 and 1/6 of 360 is 60.... i'm not sure this is linked but well it's what i can think of :/ there should be something my brain did... i hate it when i get something right and idk how i got it.... sigh

OpenStudy (anonymous):

as long as it is right thats al i care about haha thanks

OpenStudy (sasogeek):

well you should be concerned how the right answer was arrived at. if you should meet a similar question next time, i may not be there to help you

OpenStudy (anonymous):

May be she is better off guessing? :P

OpenStudy (sasogeek):

ffm... -_-

OpenStudy (anonymous):

haha saso :)

OpenStudy (anonymous):

Did you guys get it?

OpenStudy (anonymous):

anybody here?

OpenStudy (sasogeek):

no... :/

OpenStudy (anonymous):

lol, after I give you the answer, you guys will laugh :D

OpenStudy (anonymous):

I think Myin got it, she is just being silent

OpenStudy (sasogeek):

well...

OpenStudy (anonymous):

well ?

OpenStudy (sasogeek):

how is it 60...?

myininaya (myininaya):

how about you guys assume they are right triangles \[A=\frac{1}{2}bh=2160 => \frac{1}{2}(6b_1)(6h_1)=2160 => \frac{1}{2}b_1h_1=60\]

OpenStudy (anonymous):

area of the triangle is \( \frac12 \times \)base \(\times \) altitude

myininaya (myininaya):

or i mean you don't have to assume right triangles lol

OpenStudy (anonymous):

Lol, I knew it :D this is really easy as pie :D

myininaya (myininaya):

more like cake

OpenStudy (anonymous):

lol yeah chocolate cake :D

OpenStudy (sasogeek):

so how come i tend to get the answer right yet idk how i even got there.... i know i didn't guess for sure... but this is weird

OpenStudy (anonymous):

go answer my new question :)